OpenStudy (anonymous):

Suppose the area of a circle is 16π. What is its circumference?

OpenStudy (anonymous):

C=14.18

OpenStudy (anonymous):

see... are of circle=pi*r^2 => 16*pi=pi*r^2 =>16=r^2 =>r=4 now circumfrence of a circle=2*pi*r => C=2*pi*4=8*pi

OpenStudy (anonymous):

do not givee direct ans...since this is a study site not answering...learn asker how you get it..!!! @craigmo22

OpenStudy (anonymous):

K

OpenStudy (anonymous):

@zaibali.qasmi